# ORCID: https://orcid.org/0000-0001-5105-1463

### 28 Feb 2024 v1

@@ Abstract
The ongoing controversy surrounding transformative agreements, which aim to transition journal publishing to full open access, highlight the need for large-scale studies assessing the uptake of open access in hybrid journals. This includes evaluating the extent to which transformative agreements enabled open access. By combining publicly available data from various sources, including cOAlition S Journal Checker, Crossref, and OpenAlex, this study presents a novel approach that analyses over 700 agreements and nine million journal articles published in more than 11.000 hybrid journals. Estimates suggest a strong growth in open access between 2018 and 2022 from 4.3% to 15%. In 2022, 58% of hybrid open access was enabled by transformative agreements. This trend was largely driven by the three commercial publishers Elsevier, Springer Nature, and Wiley, but the open access uptake varied substantially across journals, publishers, disciplines, and country affiliations. In particular, comparing the developments in the OECD and BRICS areas revealed different publication trends relative to hybrid open access. In conclusion, estimates suggest that current levels of implementation of transformative agreements is insufficient to bring about a large-scale transition to full open access.
@@ 1 Introduction
For over two decades, hybrid open access journal publishing, which makes some articles openly available while others remain behind a paywall, has been discussed as a means for transitioning the subscription system to full open access (Prosser, 2003). The idea was that when journals increasingly publish open access articles, they could reduce revenues from subscriptions, while libraries and funders could change their funding models and shift expenditures from subscription to open access. However, initial approaches, mainly based on publication fees, also called article processing charges (APCs), did not contribute substantially to a large open access uptake. In 2009, Springer reported open access to 1% of articles in hybrid journals (Dallmeier-Tiessen et al., 2010). Other studies have also recorded low uptake. In 2011, only 1-2% of articles were open access (Björk, 2012), increasing to around 4% between 2011 and 2013 (Laakso & Björk, 2016).
@@
By the end of 2023, many transformative agreements had been implemented, but the interim outcomes were mixed. The ESAC Transformative Agreement Registry [^2] https://esac-initiative.org/about/transformative-agreements/agreement-registry/, the largest source of disclosure, recorded more than 800 transformative agreements, resulting in up to 900.000 open access articles published in both fully open access and hybrid journals according to the accompanying ESAC Market Watch [^3] https://esac-initiative.org/market-watch/. Library consortia reported increased open access volume, streamlined payment and monitoring procedures, as well as extensive utilisation of open access options by the researchers they serve (Marques & Stone, 2020; Parmhed & Säll, 2023; Pinhasi et al., 2020). The ongoing standardisation of transformative agreements has contributed to improved transparency in terms of contracts and publisher-provided article metadata (Marques et al., 2019; Pinhasi et al., 2021). However, with the growing trend toward transformative agreements, continued reliance on big deals is perceived as problematic, because it perpetuates market concentration (Butler et al., 2023; Shu & Larivière, 2023). Whether transformative agreements lead to reduced pricing remains uncertain (Borrego, 2023) and a substantial transition of hybrid journals towards full open access could not be observed (Matthias et al., 2019; Momeni et al., 2021). The focus on large commercial publishers might also increase inequality (Ross-Hellauer et al., 2022), because transformative agreements’ focus on pay to publish mainly targets institutions from high-income countries, furthering a questionable journal prestige culture (Babini et al., 2022). Besides, an editorial-board resignation raised concerns that transformative agreements might encourage publishers to maximize journal publication volume “without regard to quality” (Rasmussen, 2023).
@@
The controversies surrounding hybrid open access and transformative agreements have led to varying policy conclusions. For instance, the Association of Swedish Higher Education Institutions (Sveriges universitets- och högskoleförbund, SUHF) recommended only supporting agreements for publishing in fully open access journals  [^4] https://www.su.se/english/news/open-access-need-to-move-away-from-transformative-agreements-1.683787. Likewise, most cOAlition S funders will discontinue financial support for transformative agreements by the end of 2024 (Liverpool, 2023). The consortium also removed the majority of hybrid journals from its Transformative Journal program in 2023 due to publishers’ failure to meet self-defined open access growth targets (Brainard, 2023). By contrast, the German DEAL consortium announced a five-year transformative agreement with Elsevier starting in 2024 and also renewed its contracts with Springer Nature and Wiley until the end of 2028. Similarly, the Colombia Consortium signed the first transformative agreements in Latin America (Muñoz-Vélez et al., 2024).
@@
Despite these controversies around transformative agreements as a means of transitioning journal publishing to full open access, there is limited evidence available on the uptake of open access in hybrid journals, and the extent to which this can be attributed to transformative agreements. Previous studies have focused on specific countries (Haucap et al., 2021; Huang et al., 2020; Pölönen et al., 2020; Taubert et al., 2023; Wenaas, 2022) or publisher portfolios (Bakker et al., 2024; Fraser et al., 2023; Jahn et al., 2022; Momeni et al., 2023; Pieper & Broschinski, 2018; Schmal, 2024), while large-scale studies relied on self-reported agreement data (Moskovkin et al., 2022), or used APC pricing lists (Shu & Larivière, 2023). In particular, data availability is a limiting factor when studying the impact of transformative agreements (Bakker et al., 2024), because bibliometric databases, even though many allow the retrieval of open access articles in hybrid journals, do not directly attribute them to specific transformative agreements. Likewise, article-level open access invoicing data are only partly available (Jahn et al., 2022).
@@ 2.1 Transformative agreement and hybrid journal data
A limitation of using the Journal Checker Tool and its underlying publicly available data dump to study the development of transformative agreements over time is that expired transformative agreements are constantly removed. To address this, four different snapshots were safeguarded and combined for this study: self-archived versions from July 2021, July 2022, and May 2023, as well as the most current dump downloaded on 11 December 2023. This ensured that transformative agreements, which ended from 2021 onwards, were included, representing the majority of transformative agreements. Overall, the four combined Transformative Agreement Data dumps used in this study contained 729 of 869 agreements listed in the ESAC registry by December 2023.
The Transformative Agreement Data dumps link agreements to journals represented by journal names and the ISSN. After mapping ISSN variants to the corresponding linking ISSN (ISSN-L), as provided by the ISSN International Centre, journals were associated with publishers using the ESAC ID, a unique identifier for transformative agreements in the ESAC Transformative Agreement Registry. Furthermore, journal subjects according to the All Science Journal Classification (ASJC) were added from the Scopus journal source list as of August 2023.
@@ 2.2 Article and author metadata
After identifying hybrid journals included in transformative agreements, article metadata were retrieved from the Crossref November 2023 database snapshot for the five-year period from 2018 to 2022, according to the issued date, representing the earliest known publication date. Because Crossref metadata lacked information to distinguish between original research articles, including reviews, and other types of journal content, which are often not covered by transformative agreements (Borrego et al., 2021), only articles published in regular issues indicated by non-numeric pagination were included. Furthermore, paratext recognition was applied to exclude non-scholarly journal content such as table of contents.
@@
After retrieving the article metadata, the publication volume, including open access, was calculated for each journal. To improve the identification of hybrid journals, journals with an open access proportion above 95% were excluded. This step allowed removing additional 241 fully open access journals.
Affiliation metadata about corresponding authors are crucial for the planning and evaluation of transformative agreements because they are considered responsible for arranging open access publication (Borrego et al., 2021; Geschuhn & Stone, 2017; Schimmer et al., 2015). For this study, country and institutional affiliations were retrieved from OpenAlex. However, because corresponding author affiliation was only available for 54% of the investigated articles in OpenAlex, this study focused on first authors and their affiliations instead; approximately 90% of the investigated articles had first author affiliation metadata in OpenAlex, which is a much larger proportion than previoulsy reported for the October 2022 snapshot (Zhang et al., 2024). First authors typically contribute the most to a paper and are often considered being lead authors (Larivière et al., 2016). Related studies also assumed first authors as a proxy to examine open access payments and transformative agreements (Haucap et al., 2021; Shu & Larivière, 2023; Zhang et al., 2022).
@@ 2.3 Data analysis
Throughout this mostly automated data gathering and analysis process, Tidyverse tools (Wickham et al., 2019) for the R programming language (R Core Team, 2020) were used. The resulting data are openly available through an R data package, hoaddata (Jahn, 2023). Following Marwick et al. (2018), hoaddata contains not only the datasets used in the data analysis. It also includes code used to compile the data by connecting it to the cloud-based Google Big Query data warehouse, where scholarly big data from Crossref, OpenAlex and Unpaywall were imported, using bigrquery (Wickham & Bryan, 2023). To increase computational reproducibility, data aggregation through hoaddata was automatically carried out using GitHub Actions, a continuous integration service.
@@ 3.1 Overview
Figure 3B highlights that the majority of hybrid open access was made available through transformative agreements in 2021 and 2022, contributing 58% of the total open access article volume in 2022. However, there was also a notable increase in open access provision through other means, presumably publication fees being not invoiced through transformative agreements, from 4.1% (n = 62,625) in 2018 to 6.3% (n = 105,896) in 2022.
@@ 3.2 Publishing market
However, there are notable differences between the three large publishers. Although Elsevier published the largest volume of articles (n = 2,770,826, 34%), it recorded a comparably low number of open access articles, including those that are associated with transformative agreements. In contrast, Springer Nature and Wiley provided open access to a larger proportion of their articles (13% of Springer Nature articles and 15% of Wiley articles were open access), leading to higher open access market share (23% Springer Nature resp. 23% Wiley). This difference between Elsevier on the one hand and Springer Nature and Wiley on the other can be attributed to transformative agreements, as the latter made the majority of their open access articles available through such deals (Springer Nature 57% resp. Wiley 55%).
@@ 3.4 Comparing countries
Between 2018 and 2022, high-income countries almost exclusively dominated hybrid open access publishing through transformative agreements. During this period, first authors affiliated with institutions from Organisation for Economic Co-operation and Development (OECD) member countries published 602,050 open access articles in hybrid journals, representing 81% of the investigated open access articles. This disparity between OECD nations and other countries becomes even more evident when considering open access through transformative agreements, as 310,712 of 328,957, or 94% of open access articles were associated with such agreements.
@@ 4 Discussion
The primary aim of this study was to assess the uptake of open access in hybrid journals included in transformative agreements, which were introduced as a temporal means to support the transition of subscription-based academic publishing to full open access. This study presents a novel approach based on open data, which leverages metadata on over 700 agreements and nine million journal articles to estimate the extent to which transformation agreements contribute to the transition of this journal business model. The results highlight a strong growth in open access between 2018 and 2022, driven by an increasing number of institutions that had transformative agreements in place. However, the majority of research literature published in hybrid journals in this five-year period remained behind publisher paywalls. Growth in the adoption of open access in hybrid journals, in particular through transformative agreements, can be largely attributed to three large commercial publishers – Elsevier, Springer Nature, and Wiley – but varies substantially across journals, publishers, disciplines, and country affiliations. Despite the limitations of the data, the findings indicate that the current level of implementation of transformative agreements is insufficient to bring about a large-scale transition to fully open access.
@@
Moreover, this study presents varying levels of open access uptake through transformative agreements across journals, which can be attributed to the alignment of authors’ affiliations and the availability of such agreements at their institutions. In line with previous research findings (Butler et al., 2023; Jahn et al., 2022; Wenaas, 2022), high uptake rates were observed across hybrid journals in Social Sciences and Humanities. However, it is important to emphasise that these hybrid journals do not encompass the entire field. For example, Khanna et al. (2022) found that approximately 60% of journals utilizing the open-source publishing platform Open Journal Systems (OJS) fall within the Social Sciences and Humanities. In these fields, the majority of full open access journals are so-called “Diamond OA journals”, which do not charge publication fees (Bosman, Frantsvåg, et al., 2021). Rather, this result can be more accurately attributed to the substantial proportion of authors from high-income countries who publish in these hybrid journals included in transformative agreements, particularly when journals are of local relevance, such as by belonging to national societies or regional research scope.
@@
This large-scale study provides the first empirical evidence of the influence of transformative agreements on the transition of hybrid journals to full open access. However, several limitations needs to be acknowledged. From a data perspective, estimations of open access through transformative agreements were established by linking first author affiliations with publicly available agreement data from cOAlition S, and not through invoicing data, which is usually not shared. Moreover, this study is unable to account for the various types of transformative agreements, particularly with regard to article types and caps that limit the number of open access articles covered. Furthermore, assessing the quality of the OpenAlex and Crossref data used, particularly in terms of affiliations and article types, by combining them with established bibliometric databases such as Scopus and Web of Science was beyond the scope of this analysis (Visser et al., 2021). Overall, the methodology is designed in such a way that the results underestimate rather than overestimate the adoption of open access through transformative agreements. In addition, the data used in this study are openly available, along with the code used for this analysis.
@@
This study has practical implications for research funding and libraries. Of primary concern should be the observed differences across countries, particularly the relationship between economic development and open access adoption. From a data perspective, the reporting of open access funding, including transformative agreements, is not harmonised, but often crowd-sourced from various sources. To improve the assessment of transformative agreements, libraries and publishers need to collaborate on standards and services to publicly share information about respective journal portfolios, participating institutions and open access invoicing.

### 18 Mar 2024 v2

@@ Abstract
The ongoing controversy surrounding transformative agreements, which aim to transition journal publishing to full open access, highlight the need for large-scale studies assessing the impact of these agreements on hybrid open access. By combining publicly available data from various sources, including cOAlition S Journal Checker, Crossref, and OpenAlex, this study presents a novel approach that analyses over 700 agreements. Results suggest a strong growth in open access between 2018 and 2022 from 4.3% to 15%. During this period, 11,189 hybrid journals provided open access to 742,369 out of 8,146,958 articles, representing a five-year open access proportion of 9.1%. Authors who could make use of transformative agreements at the time of publication contributed 328,957 open access articles. In 2022, 143,615 out of 249,511 open access articles in hybrid journals (58%) were enabled by transformative agreements. This trend was largely driven by the three commercial publishers Elsevier, Springer Nature, and Wiley, but the open access uptake varied substantially across journals, publishers, disciplines, and country affiliations. In particular, the OECD and BRICS areas revealed different publication trends. In conclusion, this study suggests that current levels of implementation of transformative agreements is insufficient to bring about a large-scale transition to full open access.
@@ 1 Introduction
For over two decades, hybrid open access journal publishing, which makes some articles openly available while others remain behind a paywall, has been discussed as a means for transitioning the subscription system to full open access (Prosser, 2003). The idea was that when journals increasingly publish open access articles, they could reduce revenues from subscriptions, while libraries and funders could change their funding models and shift expenditures from subscription to open access. However, initial approaches, mainly based on publication fees, also called article processing charges (APCs), did not contribute substantially to a large open access uptake. In 2009, the publisher Springer reported that 1% of its articles in hybrid journals were open access (Dallmeier-Tiessen et al., 2010). Other studies have also recorded low uptake. In 2011, only 1-2% of articles were open access (Björk, 2012), increasing to around 4% between 2011 and 2013 (Laakso & Björk, 2016).
@@
By the end of 2023, many transformative agreements had been implemented, but the interim outcomes were mixed. The ESAC Transformative Agreement Registry [^2] https://esac-initiative.org/about/transformative-agreements/agreement-registry/, the largest resource for library consortia to publicly share their agreements, recorded more than 800 transformative agreements. These agreements resulted in the publication of up to 900.000 open access articles published in both fully open access and hybrid journals, according to the accompanying ESAC Market Watch [^3] https://esac-initiative.org/market-watch/. Library consortia reported increased open access volume, streamlined payment and monitoring procedures, as well as extensive utilisation of open access options by the researchers they serve (Marques & Stone, 2020; Parmhed & Säll, 2023; Pinhasi et al., 2020). The ongoing standardisation of transformative agreements has contributed to improved transparency in terms of contracts and publisher-provided article metadata (Marques et al., 2019; Pinhasi et al., 2021). However, with the growing trend toward transformative agreements, continued reliance on big deals is perceived as problematic, because it perpetuates market concentration (Butler et al., 2023; Shu & Larivière, 2023). Whether transformative agreements lead to reduced pricing remains uncertain (Borrego, 2023) and a substantial transition of hybrid journals towards full open access could not be observed (Matthias et al., 2019; Momeni et al., 2021). The focus on large commercial publishers might also increase inequality (Ross-Hellauer et al., 2022), because transformative agreements’ focus on pay to publish mainly targets institutions from high-income countries, furthering a questionable journal prestige culture (Babini et al., 2022). Besides, an editorial-board resignation raised concerns that transformative agreements might encourage publishers to maximize journal publication volume “without regard to quality” (Rasmussen, 2023).
@@
The controversies surrounding hybrid open access and transformative agreements have led to varying policy conclusions. For instance, the Association of Swedish Higher Education Institutions (Sveriges universitets- och högskoleförbund, SUHF) recommended only supporting agreements for publishing in fully open access journals by 2026  [^4] https://www.su.se/english/news/open-access-need-to-move-away-from-transformative-agreements-1.683787. Likewise, most cOAlition S funders will discontinue financial support for transformative agreements by the end of 2024 (Liverpool, 2023). cOAlition S also removed the majority of hybrid journals from its Transformative Journal program in 2023 due to publishers’ failure to meet self-defined open access growth targets (Brainard, 2023). By contrast, the German DEAL consortium announced a five-year transformative agreement with Elsevier starting in 2024 and also renewed its contracts with Springer Nature and Wiley until the end of 2028. Similarly, the Colombia Consortium signed the first transformative agreements in Latin America (Muñoz-Vélez et al., 2024).
@@
Despite these controversies around transformative agreements as a means of transitioning journal publishing to full open access, there is limited evidence available on the uptake of open access in hybrid journals, and the extent to which this can be attributed to transformative agreements. Previous studies have focused on specific countries (Brayman et al., 2024; Haucap et al., 2021; Huang et al., 2020; Pölönen et al., 2020; Taubert et al., 2023; Wenaas, 2022) or publisher portfolios (Bakker et al., 2024; Fraser et al., 2023; Jahn et al., 2022; Momeni et al., 2023; Pieper & Broschinski, 2018; Schmal, 2024), while large-scale studies relied on self-reported agreement data (Moskovkin et al., 2022), or used APC pricing lists (Shu & Larivière, 2023). In particular, data availability is a limiting factor when studying the impact of transformative agreements (Bakker et al., 2024), because bibliometric databases, even though many allow the retrieval of open access articles in hybrid journals, do not directly attribute them to specific transformative agreements. Likewise, article-level open access invoicing and cost data are only partly available (Jahn et al., 2022; Kramer, 2024).
@@ 2.1 Transformative agreement and hybrid journal data
A limitation of using the Journal Checker Tool and its underlying publicly available data dump to study the development of transformative agreements over time is that expired transformative agreements are constantly removed. To address this, four different snapshots were safeguarded and combined for this study: self-archived versions from July 2021, July 2022, and May 2023, as well as the most current dump downloaded on 11 December 2023. This ensured that transformative agreements, which ended from 2021 onwards, were included, representing the majority of transformative agreements. Overall, the four combined Transformative Agreement Data dumps used in this study contained 729 of 869 agreements listed in the ESAC Transformative Agreement Registry by December 2023.
The Transformative Agreement Data dumps link agreements to journals represented by journal names and the ISSN. After mapping ISSN variants to the corresponding linking ISSN (ISSN-L), as provided by the ISSN International Centre, journals were associated with publishers according to the ESAC Transformative Agreement Registry. This reflects that some portfolios may include imprints. Furthermore, journal subjects according to the All Science Journal Classification (ASJC) were added from the Scopus journal source list as of August 2023.
@@ 2.2 Article and author metadata
After identifying hybrid journals included in transformative agreements, article metadata were retrieved from the Crossref November 2023 database snapshot for the five-year period from 2018 to 2022, according to the issued date, representing the earliest known publication date. Because Crossref metadata lacked information to distinguish between original research articles, including reviews, and other types of journal content, which are often not covered by transformative agreements (Borrego et al., 2021), only articles published in regular issues indicated by numeric pagination were included. Furthermore, paratext recognition was applied to exclude non-scholarly journal content such as table of contents.
@@
After retrieving the article metadata, the publication volume, including open access, was calculated for each journal. To improve the identification of hybrid journals, journals with an open access proportion above 95% were excluded. This step allowed removing additional 134 fully open access journals. Together, these journals published 8,565 articles between 2018 and 2022.
Affiliation metadata about corresponding authors are crucial for the planning and evaluation of transformative agreements because they are considered responsible for arranging open access publication (Borrego et al., 2021; Geschuhn & Stone, 2017; Schimmer et al., 2015). For this study, country and institutional affiliations were retrieved from OpenAlex. However, because corresponding author affiliation was only available for 54% of the investigated articles in OpenAlex, this study focused on first authors and their affiliations instead; approximately 90% of the investigated articles had first author affiliation metadata in OpenAlex, which is a much larger proportion than previously reported for the October 2022 snapshot (Zhang et al., 2024). First authors typically contribute the most to a paper and are often considered being lead authors (Larivière et al., 2016). Related studies also assumed first authors as a proxy to examine open access payments and transformative agreements (Haucap et al., 2021; Shu & Larivière, 2023; Zhang et al., 2022).
@@ 2.3 Data analysis
Throughout this mostly automated data gathering and analysis process, Tidyverse tools (Wickham et al., 2019) for the R programming language (R Core Team, 2020) were used. The resulting data are openly available through an R data package, hoaddata (Jahn, 2023). Following Marwick et al. (2018), hoaddata contains not only the datasets used in the data analysis. It also includes code used to compile the data by connecting it to a cloud-based Google BigQuery data warehouse, where scholarly big data from Crossref, OpenAlex and Unpaywall were made available, using bigrquery (Wickham & Bryan, 2023). To increase computational reproducibility, data aggregation through hoaddata was automatically carried out using GitHub Actions, a continuous integration service.
@@ 3.1 Overview
Figure 3B highlights that the majority of hybrid open access was made available through transformative agreements in 2021 and 2022. In 2022, 143,615 out of 249,511 open access articles were from eligible authors, representing 58%. However, there was also a notable increase in open access provision through other means, presumably publication fees being not invoiced through transformative agreements, from 4.1% (n = 62,625) in 2018 to 6.3% (n = 105,896) in 2022.
@@ 3.2 Publishing market
However, there are notable differences between the three large publishers. Although Elsevier published the largest volume of articles (n = 2,770,826, 34%), it recorded a comparably low number of open access articles, including those that are associated with transformative agreements. In contrast, Springer Nature and Wiley provided open access to a larger proportion of their articles (13% of Springer Nature articles and 15% of Wiley articles were open access), leading to higher open access market share (23% Springer Nature resp. 23% Wiley). This difference between Elsevier on the one hand and Springer Nature and Wiley on the other can be attributed to transformative agreements, as the latter made the majority of their open access articles available through such deals (57% Springer Nature resp. 55% Wiley).
@@ 3.4 Comparing countries
Between 2018 and 2022, high-income countries almost exclusively dominated hybrid open access publishing through transformative agreements. To discern socio-economic differences, these countries were grouped according to their membership in the Organisation for Economic Co-operation and Development (OECD). During 2018 and 2022, first authors affiliated with institutions from OECD member countries published 602,050 open access articles in hybrid journals, representing 81% of the investigated open access articles. This disparity between OECD nations and other countries becomes even more evident when considering open access through transformative agreements, as 310,712 of 328,957, or 94% of open access articles were associated with such agreements.
@@ 4 Discussion
The primary aim of this study was to assess the uptake of open access in hybrid journals included in transformative agreements, which were introduced as a temporal means to support the transition of subscription-based academic publishing to full open access. This study presents a novel approach based on open data, which leverages metadata on over 700 agreements and nine million journal articles to estimate the extent to which transformation agreements contribute to the transition of this journal business model. The results highlight a strong growth in open access between 2018 and 2022, driven by an increasing number transformative agreements. However, the majority of research literature published in hybrid journals in this five-year period remained behind publisher paywalls. Growth in the adoption of open access in hybrid journals, in particular through transformative agreements, can be largely attributed to three large commercial publishers – Elsevier, Springer Nature, and Wiley – but varies substantially across journals, publishers, disciplines, and country affiliations. Despite the limitations of the data, the findings indicate that the current level of implementation of transformative agreements is insufficient to bring about a large-scale transition to full open access.
@@
Moreover, this study presents varying levels of open access uptake through transformative agreements across journals, which can be attributed to the alignment of authors’ affiliations and the availability of such agreements at their institutions. In line with previous research findings (Butler et al., 2023; Jahn et al., 2022; Wenaas, 2022), high uptake rates were observed across Social Sciences and Humanities hybrid journals. However, it is important to emphasise that these hybrid journals do not encompass the entire field. For example, Khanna et al. (2022) found that approximately 60% of journals utilising the open-source publishing platform Open Journal Systems (OJS) fall within the Social Sciences and Humanities. In these fields, the majority of full open access journals are so-called “Diamond OA journals”, which do not charge publication fees (Bosman, Frantsvåg, et al., 2021). Rather, this result can be more accurately attributed to the substantial proportion of authors from high-income countries who publish in these hybrid journals included in transformative agreements, particularly when journals are of local relevance, such as by belonging to national societies or regional research scope.
@@
This large-scale study provides the first empirical evidence of the influence of transformative agreements on the transition of hybrid journals to full open access. However, several limitations need to be acknowledged. From a data perspective, estimations of open access through transformative agreements were established by linking first author affiliations with publicly available agreement data from cOAlition S, and not through invoicing data, which is usually not shared. Moreover, this study is unable to account for the various types of transformative agreements, particularly with regard to article types and caps that limit the number of open access articles covered. Furthermore, assessing the quality of the OpenAlex and Crossref data used, particularly in terms of affiliations and article types, by combining them with established bibliometric databases such as Scopus and Web of Science was beyond the scope of this analysis (Visser et al., 2021). Overall, the methodology is designed to underestimate, rather than overestimate, the adoption of open access through transformative agreements. In addition, the data used in this study are openly available, along with the code used for this analysis.
@@
This study has practical implications for research funding and libraries. Of primary concern should be the observed differences across countries, particularly the relationship between socio-economic development and open access adoption. From a data perspective, the reporting of open access funding, including transformative agreements, is not harmonised, but often crowd-sourced from various sources. To improve the assessment of transformative agreements, libraries and publishers need to collaborate on standards and services to publicly share information about respective journal portfolios, participating institutions and open access invoicing.